# MENTOR: Guiding Hierarchical Reinforcement Learning with Human Feedback and Dynamic Distance Constraint

### v1

@@ Abstract
Hierarchical reinforcement learning (HRL) provides a promising solution for complex tasks with sparse rewards of intelligent agents, which uses a hierarchical framework that divides tasks into subgoals and completes them sequentially. However, current methods struggle to find suitable subgoals for ensuring a stable learning process. Without additional guidance, it is impractical to rely solely on exploration or heuristics methods to determine subgoals in a large goal space. To address the issue, We propose a general hierarchical reinforcement learning framework incorporating human feedback and dynamic distance constraints (MENTOR). MENTOR acts as a “mentor”, incorporating human feedback into high-level policy learning, to find better subgoals. As for low-level policy, MENTOR designs a dual policy for exploration-exploitation decoupling respectively to stabilize the training. Furthermore, although humans can simply break down tasks into subgoals to guide the right learning direction, subgoals that are too difficult or too easy can still hinder downstream learning efficiency. We propose the Dynamic Distance Constraint (DDC) mechanism dynamically adjusting the space of optional subgoals. Thus MENTOR can generate subgoals matching the low-level policy learning process from easy to hard. Extensive experiments demonstrate that MENTOR uses a small amount of human feedback to achieve significant improvement in complex tasks with sparse rewards.
1 Introduction
@@
The problem of sparse reward is consistently challenging in the domain of reinforcement learning (RL) [Reddy et al., 2019, Vecerik et al., 2017], attributing to two main factors: challenging exploration, and unstable training. In recent years, several approaches have been proposed to relieve these issues, including goal-conditional reinforcement learning [Andrychowicz et al., 2017], curiosity-driven exploration [Burda et al., 2018] and hierarchical reinforcement learning [Levy et al., 2017, Nachum et al., 2018, Hutsebaut-Buysse et al., 2022].
HRL is effective for long-horizon tasks with sparse rewards, as it divides these into more manageable subgoals, easing the issues of challenging exploration and unstable training. However, there are two difficulties in the practical application of HRL. (1) Generating instructive subgoals. Several studies have focused on proposing methods to address the challenges associated with generating efficient subgoals for guiding low-level policy. These methods involve both manual design approaches [Dietterich, 2000, Parr and Russell, 1997] and automatic generation [Zhang et al., 2022, Eysenbach et al., 2018, Park et al., 2023, Hutsebaut-Buysse et al., 2022]. However, each has drawbacks: manual design is costly and challenging to solve complex tasks [Ahn et al., 2022], while automatic generation requires extensive resources to search the entire state space [Campos et al., 2020]. (2) Completing subgoals efficiently. As for low-level, hindsight relabeling [Andrychowicz et al., 2017] modifies subgoals to convert failed transitions into successes but lacks exploratory capabilities. Curiosity-driven methods like RND prevent state revisitation [Bougie and Ichise, 2022] but can destabilize training due to reward exploration bonuses, especially with sparse rewards. At the high-level, previous research primarily focused on modifying subgoals to address non-stationarity in high-level policy updates when low-level actions fail to achieve these subgoals [Levy et al., 2017, Nachum et al., 2018]. However, these adjustments mainly address non-stationarity resulting from failures without ensuring low-level success in subgoal achievement. Levy et al. [2017] suggest penalizing high-level policies, but such penalties are ineffective if the low-level always fails.
In this paper, we introduce a new framework named MENTOR, incorporating human feedback into high-level policy learning to find better subgoals. To guide an agent in exploration, MENTOR incorporates human guidance for subgoal generation at the high-level. While humans may not accomplish complex tasks, they can give directions. Reinforcement Learning from Human Feedback (RLHF) is a technique that learns human preferences through binary comparisons of subgoals, offering directional guidance. Therefore, the high-level can receive instruction, resulting in dual advantages between HRL and RLHF.
To enable an agent to quickly complete subgoals, one low-level policy struggles to accomplish subgoals while introducing curiosity-driven exploration. Hence, we introduce Exploration-Exploitation Decoupling (EED), which uses one policy to explore while the other policy learns from the experience of the exploring policy to stabilize the training. As for the high-level, humans can provide direction, yet subgoals might not align with low-level capabilities because they may be too difficult or too easy. For cooperation between the high-level and low-level, a method called Dynamic Distance Constraint is proposed. It uses distance as a measure of subgoal difficulty and restricts the high-level subgoal space to align subgoal difficulty with low-level capabilities. We summarize the main contributions as follows:
  • $\bullet$
@@
2 related work
2.1 Hierarchical reinforcement learning
In the field of HRL, identifying meaningful subgoals within long-horizon tasks has been extensive research. This includes studies on options [Bacon et al., 2017, Harb et al., 2018, Bagaria and Konidaris, 2019], goals [Andrychowicz et al., 2017, Schaul et al., 2015, Levy et al., 2017] and skills [Eysenbach et al., 2018, Sharma et al., 2019, Park et al., 2023]. Manual-designed subgoals are costly and challenging for complex tasks [Ahn et al., 2022]. Automatic learning of meaningful subgoals without any guidance from an external expert is a significant challenge [Jinnai et al., 2019]. CSD [Park et al., 2023] aims to discover latent skills via mutual-information maximization. However, combining meaningful skills into task completion is a continuous challenge. HAC [Levy et al., 2017] addresses the challenges of sparse reward and non-stationarity at high-levels through the utilization of hindsight action transitions and hindsight goal transitions. HIRO [Nachum et al., 2018] employs a model to generate a new high-level action to rectify the high-level transitions. However, Both HAC and HIRO focus on solving non-stationary problems but fall short in task decomposition. Our research incorporates human guidance and DDC to identify appropriate task decomposition. Human insights provide subgoal direction; DDC limits subgoal difficulty.
2.2 Reinforcement Learning from Human Feedback
The surge in popularity of ChatGPT [Brown et al., 2020, Ouyang et al., 2022] has significantly boosted the recognition of RLHF in recent times. RLHF is an effective technique that can expedite AI training and bolster AI capabilities through the utilization of human feedback [Lee et al., 2021a, b, Park et al., 2022]. RLHF learns reward functions from pairwise comparison and ranking based on human preference [Lee et al., 2021a, Shah et al., 2016, Ouyang et al., 2022, Saha and Krishnamurthy, 2022]. Human intuition and experience can be incorporated as guidance in high-level decision-making, particularly in setting subgoals within the HRL framework [Bougie and Ichise, 2022]. Nevertheless, humans may struggle to offer immediate guidance that corresponds with the agent’s capabilities.
@@ 2.2 Reinforcement Learning from Human Feedback
3 Preliminary
3.1 Problem Setting
In this work, we consider high-level actions as goal states and provide rewards to the low-level policy for executing actions that result in resembling the desired goal. We define Markov Decision Process with a set of goals, characterized by the tuple $\langle{\mathcal{S}},{\mathcal{G}},{\mathcal{A}},{P},{R},\gamma\rangle$, in which $\mathcal{S}$ is the set of states; $\mathcal{G}$ is a set of goals, which is equivalent to the set $\mathcal{S}$. $\mathcal{A}$ is the set of actions; ${P(s^{{\prime}}|s,a)}$ is the transition probability function; $R(r|s^{{\prime}},g)$ is the reward function; $\gamma$ is the discount rate $\in[0,1)$. We can define the reward function $r_{g}(s)=\mathbbm{1}(\|g-s\|_{2}<\epsilon)$, which $\epsilon$ is a distance threshold determining contact. The problem of goal-conditioned task can be defined mathematically as $\pi\leftarrow\mathop{\arg\max}\limits_{\pi}\mathbb{E}_{\mathcal{T}\sim\pi(s,g)}[\sum_{t=1}^{H}\lambda^{t}r_{g}(s_{t})]$.
3.2 Hindsight Relabelling
Hindsight relabelling Andrychowicz et al. [2017] addresses sparse reward in GCRL by redefining failed transitions as successful transitions, thus generating positive rewards. Specifically, for a failed transition $(s_{t},a_{t},r_{t},s_{t+1},g)$ with $r_{t}=\mathbbm{1}(\|g_{t}-s_{t+1}\|_{2}<\epsilon)=0$, it resets the transition as $(s_{t},a_{t},1,s_{t+1},g^{\prime})$ with goal $g^{\prime}=s_{t+1}$. By utilizing this new transition, it becomes possible to train an off-policy RL algorithm with more positive rewards.
3.3 Curiosity-driven exploration
Intrinsic motivation has been utilized to encourage agents to learn about their surroundings, even when extrinsic rewards are scarce [Bellemare et al., 2016, Pathak et al., 2017, Burda et al., 2018]. In this framework, RND Burda et al. [2018] serves as the intrinsic reward. RND consists of two neural networks, represented as $f:\mathcal{S}\rightarrow{\mathbb{R}^{k}}$ and $\hat{f}:\mathcal{S}\rightarrow{\mathbb{R}^{k}}$, which are both randomly initialized and capable of transforming observations into embeddings. By keeping one network fixed and training the other network to minimize prediction distance, the intrinsic reward is formulated as $r_{rnd}(s_{t})=||\hat{f}(s_{t})-f(s_{t})||^{2}$. This intrinsic reward encourages the agent to visit novel states, as it will be higher in regions of the state space that the predictor network finds difficult to approximate, thus fostering exploration in the learning process.
4 METHODOLOGY
As shown in Figure 1, we propose MENTOR, utilizing a small amount of noisy human feedback to address the challenge of generating complex subgoals at the high-level. At the low-level, it decouples exploration and exploitation, solving the instability associated with curiosity-driven exploration. Additionally, we introduce DDC, a technique to constrain the range of high-level subgoals, aligning their difficulty with the low-level capabilities. MENTOR enables the high-level policy to set subgoals, while the low-level efficiently achieves them, facilitating rapid and stable learning. For detailed implementation, please see Appendix A.
4.1 Framework formalization
We briefly review here to introduce notation. The framework consists of four neural networks: high-level policy $\pi_{hi}:\mathcal{S}\times\mathcal{G}\rightarrow\mathcal{G}$, low-level policy $\pi_{lo}:\mathcal{S}\times\mathcal{G}\rightarrow\mathcal{A}$, reward model learned by human feedback $r_{hf}:\mathcal{S}\times\mathcal{G}\times\mathcal{G}\rightarrow[0,1]$ and RND model $f_{rnd}:\mathcal{S}\rightarrow[0,1]$ with parameters $\theta_{hi}$, $\theta_{lo}$ , $\theta_{hf}$ and $\theta_{rnd}$ respectively. In the following, we will explain how the high-level incorporates human guidance and the low-level stabilizes the training process.
@@ 4.1 Framework formalization
High-level Policy. Subgoal generation poses a significant challenge, with manual setup being costly and difficult, while automatic methods demand extensive computational resources to explore the state space. High-level subgoal generation requires macro, abstract, and generalized guidance, closely related to human preferences. RLHF offers human preferences through sample comparisons. Applied to HRL, it addresses the challenge of defining subgoals, while HRL mitigates the issue of vague and noisy preferences in RLHF, creating a mutual benefit for both HRL and RLHF.

### v2

@@ Abstract
Hierarchical reinforcement learning (HRL) provides a promising solution for complex tasks with sparse rewards of agents, which uses a hierarchical framework that divides tasks into subgoals and completes them sequentially. However, current methods struggle to find suitable subgoals for ensuring a stable learning process. To address the issue, we propose a general hierarchical reinforcement learning framework incorporating human feedback and dynamic distance constraints, termed MENTOR, which acts as a “mentor”. Specifically, human feedback is incorporated into high-level policy learning to find better subgoals. Furthermore, we propose the Dynamic Distance Constraint (DDC) mechanism dynamically adjusting the space of optional subgoals, such that MENTOR can generate subgoals matching the low-level policy learning process from easy to hard. As a result, the learning efficiency can be improved. As for low-level policy, a dual policy is designed for exploration-exploitation decoupling to stabilize the training process. Extensive experiments demonstrate that MENTOR uses a small amount of human feedback to achieve significant improvement in complex tasks with sparse rewards. Further details and code implementations can be found at https://github.com/nidesuipao/MENTOR.git.
Index Terms:
Hierarchical reinforcement learning, reinforcement learning from human feedback, dynamic distance constraint.
I Introduction
@@
The problem of sparse reward is consistently challenging in the domain of reinforcement learning (RL) [1, 2, 3], attributing to two main factors: challenging exploration, and unstable training. In recent years, several approaches have been proposed to relieve these issues, including goal-conditional reinforcement learning [4], curiosity-driven exploration [5, 6] and hierarchical reinforcement learning (HRL) [7, 8, 9, 10].
Hierarchical Reinforcement Learning (HRL) is effective for long-horizon tasks with sparse rewards, as it decomposes tasks into more manageable subgoals, mitigating challenges related to exploration and unstable training. However, there are two primary challenges in its practical applications. 1) Generating effective subgoals. To create efficient subgoals that guide low-level policies, manual design [11, 12] and automatic generation methods [13, 14, 15, 9] have been proposed. However, manual design is resource-intensive and struggles with complex tasks [16], while automatic generation demands significant computational resources to explore the entire state space [17]. 2) Efficient subgoal completion. During low-level learning, hindsight relabeling [4] adjusts subgoals to turn failed transitions into successes but lacks the capacity for effective exploration. Curiosity-driven techniques, such as Random Network Distillation (RND), prevent revisiting states [18] but can destabilize training due to reward bonuses associated with exploration, particularly in sparse reward settings. Frequent failures in low-level subgoal achievement can lead to non-stationarity at the high level. While some studies [7, 8] address these issues through hindsight mechanisms, e.g., penalizing high-level policies [7], they do not fully resolve the problem of ensuring that low-level policies consistently accomplish tasks.
I-A Related work
I-A1 Hierarchical reinforcement learning
In the field of HRL, identifying meaningful subgoals within long-horizon tasks has been extensive research. This includes studies on options [19, 20, 21], goals [4, 7, 22] and skills [14, 15, 23, 24]. Manual-designed subgoals are costly and challenging for complex tasks [16]. Automatic learning of meaningful subgoals without any guidance from an external expert is a significant challenge [25]. CSD [15] aims to discover latent skills via mutual-information maximization. However, combining meaningful skills into task completion is a continuous challenge. Director [26] introduces a practical method for learning hierarchical behaviors directly from pixels within a learned world model. [27] proposes a skill-based hierarchical reinforcement learning (SHRL) framework for solving the problem of visual navigation of a target. The SHRL framework consists of a high-level policy and three low-level skills: search, adjustment, and exploration. [28] introduces a novel framework called HIGL, which effectively reduces the action space of high-level policies by sampling informative landmarks for exploration and training the high-level policy to generate subgoals towards selected landmarks. HAC [7] addresses the challenges of sparse reward and non-stationarity at high-levels through the utilization of hindsight action transitions and hindsight goal transitions. HIRO [8] employs a model to generate a new high-level action to rectify the high-level transitions. AGILE [29] addresses the non-stationarity issue by using adversarial learning to guide the generation of compatible subgoals for the low-level policy. Previous work has centered on resolving non-stationary issues or on task decomposition strategies, yet often faces the challenge of sparse rewards, prompting unnecessary exploration. Our study introduces human guidance and DDC to decompose tasks effectively. Human guidance directs subgoals, reducing reward sparsity, while DDC controls subgoal difficulty, synchronizes learning across levels, and relieves non-stationarity issues.
I-A2 Reinforcement Learning from Human Feedback
The surge in popularity of ChatGPT [30, 31] has significantly boosted the recognition of RLHF in recent times. RLHF is a technique for aligning the functionalities of Large Language Models (LLMs) with human ethical considerations and preference, achieved by integrating human feedback into the learning process [32, 33, 34, 35, 36]. For instance, [37] demonstrates how human preferences can be integrated into text-to-image generation models, aligning them with human aesthetics. [38] introduces a GAN-augmented reinforcement learning strategy that efficiently learns robot behaviors through human preferences, significantly reducing the reliance on human demonstrations. RLHF learns reward functions from pairwise comparison and ranking based on human preference [31, 32, 39, 40]. Human intuition and experience can be incorporated as guidance in high-level decision-making, particularly in setting subgoals within the HRL framework [18]. Nevertheless, humans may struggle to offer immediate guidance that corresponds with the agent’s capabilities.
I-B Our Contribution
We introduce a novel framework, MENTOR, which integrates human guidance into the high-level policy learning process. This is achieved through RLHF, a method that trains a reward model by learning human preferences via binary comparisons of subgoals. MENTOR utilizes this reward model to generate subgoals at the high-level, effectively steering the agent towards optimal behavior. Additionally, we introduce DDC. It measures subgoal difficulty by distance and adjusts the subgoal space accordingly. To enable an agent to quickly complete subgoals at the low-level, we introduce Exploration-Exploitation Decoupling (EED), which uses one policy to explore while the other policy learns from the experience of the exploring policy to stabilize the training. We summarize the main contributions as follows:
@@ I-B Our Contribution
II Preliminary
II-A Problem Setting
Define Markov Decision Process with a set of goals, characterized by the tuple $\langle{\mathcal{S}},{\mathcal{G}},{\mathcal{A}},{P},{r},\gamma\rangle$, where $\mathcal{S}$, $\mathcal{G}$, and $\mathcal{A}$ are the sets of state, goal, and action, respectively, ${P(s^{{\prime}}|s,a)}$ is the transition probability function, $r_{\text{g}}(s,g):\mathcal{S}\times\mathcal{G}\to\mathbb{R}$ is the reward function, and $\gamma$ is the discount rate $\in[0,1)$. The objective is to find an optimal policy $\pi^{*}$ such that
$\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\mathcal{T}\sim\pi}\left[\sum_{t=1}^{H}\gamma^{t}r_{\text{g}}(s_{t},g_{t})\right],$
where $\mathcal{T}$ denotes a trajectory generated under the policy $\pi$ starting from an initial state. The reward function can be defined as $r_{\text{g}}(s,g)=\mathbbm{1}(\|g-s\|_{2}<\epsilon)$, which $\epsilon$ is a distance threshold determining contact.
II-B Hindsight Relabelling
Hindsight relabelling [4] addresses sparse reward in GCRL by redefining failed transitions as successful transitions, thus generating positive rewards. Specifically, for a failed transition $(s_{t},a_{t},r_{t},s_{t+1},g)$ with $r_{t}=r_{\text{g}}(s_{t+1},g)=0$, it resets the transition as $(s_{t},a_{t},1,s_{t+1},g^{\prime})$ with goal $g^{\prime}=s_{t+1}$. By utilizing this new transition, it becomes possible to train an off-policy RL algorithm with more positive rewards. [4] also delve three heuristic methods for goal relabeling to enhance learning efficiency: (1) Future Sampling, which selects goals from subsequent states within the same trajectory; (2) Episode-Specific Random Sampling, which randomly selects goals from the same episode without considering their order; (3) Random Sampling, which chooses new goals from the entire dataset. In our implementation, we have opted for the Future Sampling strategy.
II-C Curiosity-driven Exploration
Intrinsic motivation has been utilized to encourage agents to learn about their surroundings, even when extrinsic rewards are scarce [5, 41, 42]. Curiosity-driven exploration is an intrinsic motivation strategy, exemplified by algorithms like RND, which fosters learning through the agent’s desire to discover new information. In this framework, RND [5] serves as the intrinsic reward. RND consists of two neural networks, represented as $f:\mathcal{S}\rightarrow{\mathbb{R}^{k}}$ and $\hat{f}:\mathcal{S}\rightarrow{\mathbb{R}^{k}}$, which are both randomly initialized and capable of transforming observations into embeddings. By fixing one network and training the other to minimize the prediction error, we follow the distillation optimization process:
$\mathbb{E}_{s\sim\rho_{\pi(s)}}\min_{\hat{f}}\|\hat{f}(s)-f(s)\|^{2},$ (1)
where $s$ is sampled from the replay buffer. Upon the agent’s interaction with the environment, yielding the current state $s_{t}$, we proceed to compute RND reward:
$r_{\text{rnd}}(s_{t})=\|\hat{f}(s_{t})-f(s_{t})\|^{2}.$ (2)
This RND reward encourages the agent to visit novel states, as it will be higher in regions of the state space that the predictor network finds difficult to approximate, thus fostering exploration in the learning process. To ensure effective policy optimization and stability in reinforcement learning, it is essential that the RND reward, as an intrinsic reward, is normalized to align with the scale of extrinsic rewards:
$r_{\text{intri}}(s_{t})=\frac{r_{\text{rnd}}(s_{t})-\mu(r_{\text{rnd}})}{\sigma(r_{\text{rnd}})}.$ (3)
We determine the normalized intrinsic reward for state $s_{t}$ by converting the RND reward into a Z-score, aligning it with the mean $\mu(r_{\text{rnd}})$ and scaling it according to the standard deviation $\sigma(r_{\text{rnd}})$.
III Methodology
@@
HRL consists of multiple levels, each focused on solving specific problems. The high-level is responsible for providing overall guidance, while the low-level handles task execution. This difference in responsibilities between the levels requires distinct reward function characteristics. Depending on unique features of diverse levels, we propose MENTOR, shown in Figure 1(c). At the high-level, it utilizes RLHF and DDC to address the challenge of generating instructive subgoals shown in Figure 1(a). At the low-level, it decouples exploration and exploitation, solving the instability associated with curiosity-driven exploration shown in Figure 1(b).
Before introducing the framework, we briefly review here to introduce notation. The framework consists of four neural networks: high-level policy $\pi_{\mathrm{hi}}:\mathcal{S}\times\mathcal{G}\rightarrow\mathcal{G}$, low-level policy $\pi_{\mathrm{lo}}:\mathcal{S}\times\mathcal{G}\rightarrow\mathcal{A}$, reward model learned by human feedback $r_{\text{hf}}:\mathcal{S}\times\mathcal{G}\times\mathcal{G}\rightarrow[0,1]$, RND model $f_{\text{rnd}}:\mathcal{S}\rightarrow[0,1]$ and distance model $d:\mathcal{G}\times\mathcal{G}\rightarrow[0,1]$ with parameters $\theta_{\mathrm{hi}}$, $\theta_{\mathrm{lo}}$ , $\theta_{\text{hf}}$, $\theta_{\text{rnd}}$ and $\theta_{d}$ respectively. The distance model, denoted as $d$, operates within the Dynamic Distance Constraint. In a given episode, $\pi_{\mathrm{hi}}$ initially proposes subgoal $g^{\text{sub}}_{0}$ based on current state $s$ and environment goal $g^{\text{env}}$, followed by task given by $\pi_{\mathrm{lo}}$ to achieve $g^{\text{sub}}_{0}$. If $\pi_{\mathrm{lo}}$ succeeds and the episode remains active, $\pi_{\mathrm{hi}}$ subsequently issues a new subgoal $g^{\text{sub}}_{1}$. Thus, the high-level trajectory is represented as ($s_{t_{0}},s_{t_{1}},...,s_{t_{n-1}}$), where ${t_{i}}$ signifies the moment of the $i$-th subgoal generation by $\pi_{\mathrm{hi}}$. Concurrently, the low-level trajectory unfolds as ($s_{0},s_{1},...,s_{L-1}$), $L$ denotes the trajectory length.
III-A RLHF and Dynamic Distance Constraint in High-level
III-A1 Subgoal generation using RLHF
@@
Subgoal generation poses a significant challenge, with manual setup being costly and difficult, while automatic methods demand extensive computational resources to explore the state space. High-level subgoal generation requires macro, abstract, and generalized guidance, closely related to human preferences. RLHF offers human preferences through sample comparisons. Despite the inherent uncertainty and noise in the human feedback acquired through pairwise preference comparisons, this approach for reward model learning effectively guides the high-level policy learning because HRL mitigates the issue of uncertainty and noisy preferences in RLHF, creating a mutual benefit for both HRL and RLHF.